# Perspectives for hyperon and hypernuclei physics

### v1

@@ II.7 Theoretical progress on hypernuclear studies
For double hypernuclei, there were several data points by KEK-E176-E373: the separation energy ( of He was MeV Hiyama and Nakazawa (2018) and the observed of the excited state of Be was MeV Hiyama and Nakazawa (2018). In addition, in 2009, Be was observed and the observed was MeV. Afterwards, the search for the double hypernuclei by J-PARC E07 was performed at J-PARC. By this experiment, of Be was observed and the observed was MeV. It should be noted that it is difficult to identify states for the observed s experimentally. Therefore, it is highly desirable to interpret the states theoretically. One successful example is Be. In 2009, new data of double was observed as the Hida event Hiyama and Nakazawa (2018) as mentioned above and there are two possibilities for interpretation: One is Be with MeV, and the other is Be with MeV. Furthermore, it is uncertain whether this is an observation of a ground state or an excited state. To interpret this event, one of the present authors (E.H.) performed a study of the five-body cluster model by assuming that this event was an observation of Be Hiyama and Nakazawa (2018). In this calculation, it is essential to reproduce the observed properties of all subcluster systems composed of two, three, or four constituent particles. In this calculation, the interactions are adjusted to reproduce energies of all subcluster systems, such as , , , , , systems. The energies of the ground state and the excited state of Be are in good agreement with the observed data with no adjustable parameter. The calculated energy level is shown in Fig. 3.
@@
It is also important to search for a bound hypernucleus for the study of the interaction. Experimentally, a bound C(N) hypernucleus was observed by emulsion Nakazawa et al. (2015); Hiyama and Nakazawa (2018); Hayakawa et al. (2021); Yoshimoto et al. (2021). By this experimental data, one found that the interaction should be attractive. In addition, the latest femtoscopic data from collisions at LHC Acharya et al. (2019) indicate that the spin-isospin averaged interaction is attractive at low energies.
@@
For the system, we have two total isospins, and 1. For the system, the use of the ESC08c potential leads to bound states for and , while there is no bound state using the HAL QCD potential. For , both the ESC08c and HAL QCD potentials lead to a bound state, although the energies are dependent on the interaction employed.
@@
In the sector, one has started to obtain information on the and interactions. It is essential to obtain further information in this sector, for instance, the -wave interaction, the spin-isospin terms of the interaction, etc. One should wait for future experiments.
@@ III Summary and outlook
Recent advances in hypernuclear physics have provided profound insights into the strangeness in nuclear systems, driven by experimental techniques and theoretical innovations. Light hypernuclei, such as the hypertriton (H), are pivotal benchmarks for probing the -nucleon interaction. Notably, discrepancies in the binding energy of hypertriton between STAR and ALICE underscore the need for systematic studies to resolve these uncertainties. Future experiments at FAIR, J-PARC, and upgraded LHC facilities aim to deliver high-statistics data to clarify these inconsistencies and explore neutron-rich hypernuclei.
@@
Charge symmetry breaking in the hypernuclear system has been an unsolved puzzle on both the experimental and theoretical sides for a considerable time. The most recent experimental measurements from the STAR Collaboration of H and He in Au+Au collisions at =3 GeV showed that the CSB effect in their excited states have a similar magnitude but opposite sign as their ground states. However, this measurement suffered from large uncertainties and could not draw any solid conclusion about the CSB puzzle. The STAR Collaboration has taken significantly larger statistics at = 3 GeV. The analysis is ongoing, as one should expect to suppress the uncertainties in this new data. Measurements with electron and kaon beams are also in progress at MAMI-A1 Shao et al. (2025), JLab, and J-PARC. As demonstrated by the measurements of the STAR Collaboration, heavy-ion collisions at several GeV are excellent platforms for generating large statistics of light hypernuclei. Future experiments with heavy-ion beams, such as the High-Intensity Heavy-Ion Accelerator Facility (HIAF) Zhou and Yang (2022) and the Facility for Antiproton and Ion Research (FAIR), are also possible for studying charge symmetry breaking (CSB) in hypernuclei.
@@
In the theoretical study of hyperon-nucleon and hyperon interactions, non-relativistic chiral effective potentials have been constructed up to the next-to-next-to-leading order. In contrast, the relativistic chiral potentials are only constructed at the leading order. Going to higher orders in both the non-relativistic and the relativistic frameworks requires more precise data to fix all the relevant low-energy constants. The recent measurement of differential cross sections by the J-PARC E40 experiment Miwa et al. (2022) and the high-energy cross sections measured by CLAS Rowley et al. (2021) are of tremendous help in this perspective. In addition to moving to higher orders in the two-body sector, studies of three-body hyperon-nucleon-nucleon interactions can play a crucial role in understanding various relevant issues, such as the hyperon puzzle in neutron stars.
@@
Bare interactions provide necessary inputs in studies of hypernuclei. Conventional theoretical methods, such as the GEM method, need to be extended to treat a wider range of hadrons. Currently, it can treat five body systems. Another promising approach is the lattice effective field theory approach Le et al. (2025). Relativistic few-body calculations utilizing relativistic chiral forces are in progress and may provide unique insights into many unsolved problems. For heavy nuclei/hypernuclei out of reach of ab initio methods, one can first derive effective potentials using bare chiral hyperon/nucleon forces, and then use them as inputs with density functional theories Yang et al. (2022).

### v2

@@ II.7 Theoretical progress on hypernuclear studies
For double hypernuclei, there were several data points by KEK-E176-E373: the separation energy ( of He was MeV Hiyama and Nakazawa (2018) and the observed of the excited state of Be was MeV Hiyama and Nakazawa (2018). The theoretical aspects of the H, He, and He had been studied in the Jacobi no-core shell mode (NCSM) with chiral interactions Le et al. (2021a). In addition, in 2009, Be was observed and the measured was MeV. Afterwards, the search for the double hypernuclei by J-PARC E07 was performed at J-PARC. By this experiment, of Be was observed and the measured was MeV. It should be noted that it is difficult to identify states for the observed s experimentally. Therefore, it is highly desirable to interpret the states theoretically. One successful example is Be. In 2009, new data of double was observed as the Hida event Hiyama and Nakazawa (2018); Gal and Millener (2011) as mentioned above and there are two possibilities for interpretation: One is Be with MeV, and the other is Be with MeV. Furthermore, it is uncertain whether this is an observation of a ground state or an excited state. To interpret this event, one of the present authors (E.H.) performed a study of the five-body cluster model by assuming that this event was an observation of Be Hiyama et al. (2010); Hiyama and Nakazawa (2018). In this calculation, it is essential to reproduce the observed properties of all subcluster systems composed of two, three, or four constituent particles. In this calculation, the interactions are adjusted to reproduce energies of all subcluster systems, such as , , , , , systems. The energies of the ground state and the excited state of Be are in good agreement with the observed data with no adjustable parameter. The calculated energy level is shown in Fig. 3.
@@
It is also important to search for a bound hypernucleus for the study of the interaction. Experimentally, a bound C(N) hypernucleus was observed by emulsion Nakazawa et al. (2015); Hiyama and Nakazawa (2018); Hayakawa et al. (2021); Yoshimoto et al. (2021). By this experimental data, one found that the interaction should be attractive. In addition, the latest femtoscopic data from collisions at LHC Acharya et al. (2019a) indicate that the spin-isospin averaged interaction is attractive at low energies.
@@
For the system, we have two total isospins, and 1. For the system, the use of the ESC08c potential leads to bound states for and , while there is no bound state using the HAL QCD potential. For , both the ESC08c and HAL QCD potentials lead to a bound state, although the energies are dependent on the interaction employed. Meanwhile, three shallow bound states for the system with , and are predicted in the Jacobi NCSM with the NLO(500) chiral interactions Le et al. (2021b). It should be noted that some parameters of the chiral interactions used in Ref. Le et al. (2021b) have been retuned to the results of the original chiral baryon-baryon interaction, which was necessary because in the application to study hypernuclei the channel of the original potential had to be omitted.
@@
In the sector, one has started to obtain information on the and interactions. It is essential to obtain further information in this sector, for instance, the -wave interaction, the spin-isospin terms of the interaction, etc. New experiments on hypernuclei, scattering experiments, and correlation functions together with improved theoretical calculations will help to shed further light on the interactions.
@@ III Summary and outlook
Recent advances in hypernuclear physics have provided profound insights into the strangeness in nuclear systems, driven by experimental techniques and theoretical innovations. The measured properties of hypernuclei have been collected and evaluated in the online database (https://hypernuclei.kph.uni-mainz.de/). Light hypernuclei, such as the hypertriton (H), are pivotal benchmarks for probing the -nucleon interaction. Notably, discrepancies in the binding energy of hypertriton between STAR and ALICE underscore the need for systematic studies to resolve these uncertainties. Future experiments at FAIR, J-PARC, and upgraded LHC facilities aim to deliver high-statistics data to clarify these inconsistencies and explore neutron-rich hypernuclei.
@@
Charge symmetry breaking in the hypernuclear system has been an unsolved puzzle on both the experimental and theoretical sides for a considerable time. The most recent experimental measurements from the STAR Collaboration of H and He in Au+Au collisions at =3 GeV showed that the CSB effect in their excited states have a similar magnitude but opposite sign as their ground states. However, this measurement suffered from large uncertainties and could not draw any solid conclusions about the CSB puzzle. The STAR Collaboration has taken significantly larger statistics at = 3 GeV. The analysis is ongoing, as one should expect to suppress the uncertainties in these new data. Measurements with electron and kaon beams are also in progress at MAMI-A1 Shao et al. (2025), JLab, and J-PARC. As demonstrated by the measurements of the STAR Collaboration, heavy-ion collisions at several GeV are excellent platforms for generating large statistics of light hypernuclei. Future experiments with heavy-ion beams, such as the High-Intensity Heavy-Ion Accelerator Facility (HIAF) Zhou and Yang (2022) and the Facility for Antiproton and Ion Research (FAIR), are also possible for studying charge symmetry breaking in hypernuclei.
@@
In the theoretical study of hyperon-nucleon and hyperon interactions, non-relativistic chiral effective potentials have been constructed up to the next-to-next-to-leading order. In contrast, the relativistic chiral potentials are only constructed at the leading order. Going to higher orders in both the non-relativistic and the relativistic frameworks requires more precise data to fix all the relevant low-energy constants. The recent measurement of differential cross sections by the J-PARC E40 experiment Miwa et al. (2022) and the femtoscopic correlation functions Acharya et al. (2022, 2019b, 2020a) are of tremendous help in this perspective. In addition to moving to higher orders in the two-body sector, studies of three-body hyperon-nucleon-nucleon interactions can play a crucial role in understanding various relevant issues, such as the hyperon puzzle in neutron stars.
@@
Bare interactions provide necessary inputs in studies of hypernuclei. Conventional theoretical methods, such as the GEM method, need to be extended to treat a wider range of hadrons. Currently, it can treat five body systems. Another promising approach is the lattice effective field theory approach Hildenbrand et al. (2024). Relativistic few-body calculations utilizing relativistic chiral forces are in progress and may provide unique insights into many unsolved problems. For heavy nuclei/hypernuclei out of reach of ab initio methods, one can first derive effective potentials using bare chiral hyperon/nucleon forces, and then use them as inputs with density functional theories Yang et al. (2022).